OpenStudy (anonymous):

please help A box of mass 21 kg sits on an inclined surface with an angle of 10. What is the component of the weight of the box along the surface? A. 203 N B. 59 N C. 36 N D. 113 N

OpenStudy (shamim):

Weight of 21 kg=mg=21*9.8=?

OpenStudy (shamim):

The component of weight along the surface=mg*sin10=21*9.8*sin10=?

OpenStudy (shamim):

Response plz!!!!!!

OpenStudy (anonymous):

I got 203

OpenStudy (anonymous):

Is that right?

OpenStudy (shamim):

I think u r not right!!!

OpenStudy (shamim):

Can u tell me Sin10=?

OpenStudy (shamim):

U should calculate it by calculator